# RESPOND: Responsive Engagement Strategy for Predictive Orchestration and Dialogue

### 23 Mar 2026 v1

@@ 1. Introduction
This paper introduces RESPOND (Responsive Engagement Strategy for Predictive Orchestration in Dialogue), a framework that bridges rigid machine turn-taking and the dynamic qualities of human conversation. RESPOND enables agents to generate both timely backchannels and cooperative turn claims, simulating a full-duplex architecture in which the system listens and speaks concurrently. Built on streaming ASR and incremental semantics, RESPOND continuously analyzes unfolding speech, and its Predictive Orchestration Engine forecasts opportune moments for interjection, mirroring the fluidity of natural dialogue.
@@
This level of control is achieved architecturally through lightweight conditioning (e.g., feature-wise modulation), which allows for adjustments without retraining the core model (Perez et al., 2017). By making conversational behavior steerable and transparent, RESPOND aligns with intelligent user interfaces that prioritize user and designer agency.
@@ 2.1. Predictive turn-taking
2.2. Backchannel and turn claim prediction.
@@ 3.2. Model
The backbone of our system is the Qwen3-0.6B language model (Qwen et al., 2025), a 0.6B-parameter Transformer designed for efficiency. We fine-tune the model using Low-Rank Adaptation (LoRA (Hu et al., 2021)) to adapt it for turn-taking prediction while keeping the parameter footprint manageable. The model outputs a classification over three categories: turn claim, backchannel, or stay silent.
@@ 3.3. Controllability
We condition the model on these control values using Feature-wise Linear Modulation (FiLM) layers (Perez et al., 2017; Dumoulin et al., 2018). Given a hidden representation $h\in\mathbb{R}^{d}$ from the backbone encoder, we compute:
@@ 6. Limitations and Future Work
GenAI Usage Disclosure
Generative AI tools were used only for minor grammar and language refinement during the preparation of this manuscript. No generative AI tools were used for producing scientific content, designing experiments, analyzing data, writing code, or creating figures. All research contributions, results, and interpretations are the sole work of the authors.

### 30 Mar 2026 v2

@@ 1. Introduction
This paper introduces RESPOND (Responsive Engagement Strategy for Predictive Orchestration and Dialogue), a framework that bridges rigid machine turn-taking and the dynamic qualities of human conversation. RESPOND enables agents to generate both timely backchannels and cooperative turn claims, simulating a full-duplex architecture in which the system listens and speaks concurrently. Built on streaming ASR and incremental semantics, RESPOND continuously analyzes unfolding speech, and its Predictive Orchestration Engine forecasts opportune moments for interjection, mirroring the fluidity of natural dialogue.
@@
This level of control is achieved architecturally through lightweight conditioning (e.g., feature-wise modulation), which allows for adjustments without retraining the core model (Perez et al., 2018). By making conversational behavior steerable and transparent, RESPOND aligns with intelligent user interfaces that prioritize user and designer agency.
@@ 2.1. Predictive turn-taking
2.2. Backchannel and turn claim prediction
@@ 3.2. Model
The backbone of our system is the Qwen3-0.6B language model (Qwen et al., 2025), a 0.6B-parameter Transformer designed for efficiency. We fine-tune the model using Low-Rank Adaptation (LoRA (Hu et al., 2022)) to adapt it for turn-taking prediction while keeping the parameter footprint manageable. The model outputs a classification over three categories: turn claim, backchannel, or stay silent.
@@ 3.3. Controllability
We condition the model on these control values using Feature-wise Linear Modulation (FiLM) layers (Perez et al., 2018; Dumoulin et al., 2018). Given a hidden representation $h\in\mathbb{R}^{d}$ from the backbone encoder, we compute: